Systems Staff Meeting Notes 3/6/00

Systems Staff Meeting Notes
March 6, 2000

Present: Rich Bennett, Will Chaney, Bill Covey, Debra Harris, Winston Harris, Bobbie Meng, Hilton Soto, Ying Tang
Announcements
  1. Rich's mother is 85 today. He & Denise will celebrate their 8th wedding anniversary tomorrow.
  2. Hilton & Everett both ran a 15K race in Jacksonville this weekend.
Ongoing Systems Updates
Equipment Received/Installed/Repaired
  1. Lindsay is out this week, so Bill has to do the ordering.
  2. The Dell order for 24 work stations has gone out.
  3. The order for the pieces of the Terrabyte server should sail now that it's been gotten through Purchasing.
  4. The new Systems vacuum is on its way. Will use it first in the closet in room 219. The closet is in the process of being cleaned; Will took bolt cutters to the old coax.
Security
  1. There's a virus on NT2 - it's been running a check count every 10 minutes for the past 5 days. It's a Visual Basic scripting virus called 'FreeLink'. It gets trapped on the server & cannot spread, but we are also unable to determine where it's coming from. If anyone runs across any information on FreeLink, please pass it on to Bill Covey.
Liaisons
  1. Had the first administrative orientation session this morning. Everyone expected to show, did. The first technical session will be held this afternoon.
Networking
  1. Six new switches have been installed in room 219 (closet).
  2. Lana is out this week, but she has finished documenting the wiring in the closet of room 2D Smathers and the first floor of the north tower.
  3. The fiber farm switch has been installed in MSL.
  4. The converter was moved from Music to AFA. AFA now has 100M feed.
  5. Music is scheduled for rewiring this week.
  6. West wiring should start this week.
  7. Four of the new machines that have been ordered will be destined for Journalism - to replace the SNA terminals. The IPs need to have Circ access.
  8. Bobbie notes that the IPs on subnet 10 are running low. Bill will do a nose count on it, but it is difficult to expand upon. Will do some subnet pinging to see what responds.
Software Projects Status
  1. Q&A and Library Statistics are ongoing.
  2. Ying is making good progress with the web interface for the Library Statistic database.
  3. Small modifications were made to the Personnel and Travel databases.
  4. LASSIE modifications - Marie Kennedy met with Winston and Hilton about some changes. Bill cautions to run any changes by Carolyn before implementation. The interface is extremely complex - it could be more user friendly. Rich and Winston will need to meet to discuss the integration of LASSIE and IPs.
DP Coordinator
  1. The phones were not going to voice mail. The Telecom guy came by this morning and said that the voice mailbox for the building was full. He reset it, it's working now. We may need to urge staff members to remember to delete old messages in their voice mail.
Calendar
  1. Suzy is out until noon today.
  2. Everett is out all day today.
  3. Bobbie will be out March 9 & 10.
  4. Hilton will be out March 10.
  5. Rich has jury duty March 13.
  6. Debra will be out April 18-21.


Back to the Systems Homepage

Last updated March 13, 2000
by Debra Harris
debharr@mail.uflib.ufl.edu